Find the indicated term of the sequence.
a7 of 9, 45, 225, 1125 . . .

Answer by fractalier(6550)

This is a geometric sequence with a1 = 9 and common ratio r = 5...
a7 = a1 * r^(7-1) = 9*5^6 = 140625